OpenStudy (anonymous):

NEED an answer ASAP!!!!!!!! I know that you can change this 1/(x^2(x+1)) into: 1/x^2 +1/(x+1) -1/x However, I dont know how, could someone explain that to me? please

OpenStudy (anonymous):

Partial fractions.

OpenStudy (bahrom7893):

1/(x^2(x+1)) = A/x + B/x^2 + C/(x+1)

OpenStudy (bahrom7893):

MALEVOLENCE!!!!

OpenStudy (anonymous):

a blast from the past!

OpenStudy (anonymous):

ok and why is the A/x minus?

OpenStudy (anonymous):

lol?

OpenStudy (anonymous):

You have to solve for the constants.

OpenStudy (anonymous):

ohhh i see.. ok thanks lemme try it

OpenStudy (bahrom7893):

A/x + B/x^2 + C/(x+1) Ax(x+1) + B(x+1) + Cx^2 = 1 Ax^2 + Ax + Bx + B + Cx^2 = 1 Ax^2+Cx^2 + Ax+Bx + B = 1 (A+C)x^2 + (A+B)x + B = 0x^2 + 0x + 1

OpenStudy (bahrom7893):

A+C = 0 A+B = 0 B = 1 A = -1 C = -A = 1

OpenStudy (bahrom7893):

Thus: A/x + B/x^2 + C/(x+1) = (-1/x) + (1/x^2) + (1/(x+1))

OpenStudy (anonymous):

ok i got it bahrom, thanks a lot buddy :)

OpenStudy (anonymous):

i know i have done this sometimes last year, but i forgot it haha

OpenStudy (bahrom7893):

np sal :)